Suppose that X represents one of two positive numbers whose sum is 50. Determine a function F that represents the product of these two numbers. For what two such numbers is the product equal to 400?
Found 2 solutions by RAY100, stanbon:
Answer by RAY100(1637)

Let x and y be numbers
.
x+y =50,,,,or x= 50-y
.
product = x*y = 400,,,,subst
.
(50-y) * y = 400
.
50y -y^2 =400
.
0= y^2 -50y +400
.
(y-10) (y-40) = 0
.
y=10, or,,,y=40
.
whereas,, x= 40 or x=10
.
therefore,,,numbers are 10,,,and 40
.
.
check 10+40 = 50,,,ok,,,,,,10*40 = 400,,,ok
.

Answer by stanbon(75887)

Suppose that X represents one of two positive numbers whose sum is 50. Determine a function F that represents the product of these two numbers. For what two such numbers is the product equal to 400?
----------------
1st number: x
2nd number: 50-x
--------------------------
Product = 400
x(50-x)-400 = 0
-x^2+50x-400 = 0
x^2 - 50x + 400 = 0
(x-10)(x-40) = 0
x = 10 and x = 40
===============================
Cheers,
Stan H.